# Exploring the Limits of Vision-Language-Action Manipulations in Cross-task Generalization

### v2

@@ Abstract
Figure 1: The proposed AGNOSTOS benchmark evaluates cross-task zero-shot generalization through two difficulty levels. Level-1 testing involves 13 unseen tasks sharing partial similarity (objects or motions) with seen tasks. Level-2 testing has 10 unseen tasks from entirely novel scenarios, requiring stronger generalization capabilities. We systematically assess three broad categories of vision-language-action models, revealing critical limits in their ability to adapt to unseen tasks.
@@ 1 Introduction
Vision-Language-Action (VLA) models [3, 4, 5, 6, 2, 7, 8, 1, 9] have motivated a new era of robotic manipulation [10, 11, 12, 13, 14, 15, 16] by integrating visual perception, language understanding, and action generation. Through large-scale pre-training on diverse data, including human videos [17, 18, 19], real or simulated cross-embodiment robotic demonstrations [3, 20, 21], VLA models can effectively generalize across visual variations within known tasks (i.e., within-task generalization of seen tasks), such as handling objects in novel scenes or with altered properties. However, the true promise of VLAs lies in their capacity to generalize across tasks: to handle previously unseen combinations of objects, goals, and actions without prior exposure. This capability-cross-task zero-shot generalization-is essential for real-world deployment, where robots are expected to tackle novel tasks as they arise dynamically.
Despite the rapid progress in VLA research, most prior work [3, 4, 7, 8, 1] has focused on generalization testing in real-world environments. However, these evaluations are typically non-reproducible, and rarely target cross-task (i.e., unseen task) zero-shot generalization. Recently, many works [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] have been devoted to developing comprehensive simulated benchmarks, which could be utilized to evaluate the generalization of VLA models. While promising, these efforts mainly assess within-task generalization, leaving cross-task zero-shot generalization largely unexplored.
To address this critical gap, we present AGNOSTOS, a novel benchmark for evaluating cross-task zero-shot generalization in robotic manipulation. Built on RLBench [25], our benchmark comprises 23 unseen tasks that are carefully curated to differ from 18 commonly used seen training tasks [34, 35]. As shown in Figure 1, to probe different aspects of generalization, these unseen tasks are categorized into two difficulty levels:
@@
Motivated by the success of in-context learning [42, 43, 44] in large language models (LLMs), we propose a Cross-task In-context Manipulation (X-ICM) method to address the challenges of cross-task zero-shot generalization. X-ICM uses demonstrations from seen tasks as in-context examples to prompt LLMs to generate action plans for unseen tasks. A key challenge under this setup is selecting relevant demonstrations; irrelevant prompts fail to activate appropriate knowledge in the LLM, leading to poor cross-task predictions. To address this, we introduce a dynamics-guided sample selection strategy. This approach leverages learned representations of task dynamics—captured by predicting final observations from initial states and task descriptions—to identify relevant demonstrations across tasks. Guided by the learned dynamics, the dynamics-aware prompts can be formed that enable X-ICM to elicit high-quality action predictions from LLMs, even for unfamiliar, unseen tasks.
@@
  • We introduce AGNOSTOS, the first benchmark to systematically evaluate cross-task zero-shot generalization in robotic manipulation for VLAs, with 23 unseen tasks spanning two levels of generalization difficulty.
  • We propose X-ICM, a novel method that combines in-context prompting and dynamics-guided sample selection to enhance cross-task zero-shot transfer in VLA models.
@@ 2 Related Works
Benchmarks for Vison-Language-Action Models. Evaluating the generalization capabilities of VLA models requires robust and comprehensive benchmarks. While existing VLAs mainly focus on real-world testing, which suffers from reproducibility issues and rarely focuses systematically on zero-shot generalization to unseen tasks. Recently, many simulated benchmarks [22, 23, 24, 25, 26, 27, 28, 30, 31, 32, 33] have been proposed to offer a controlled environment for rigorous evaluation. Colosseum [31] enables evaluation of models across 14 axes of perturbations, including changes in color and texture, etc. GemBench [32] designs four levels of generalization, spanning novel placements, rigid and articulated objects, and complex long-horizon tasks. We find that these benchmarks have predominantly concentrated on evaluating visual variations within tasks, assessing how well models can generalize to new scenes or altered attributes of objects within the same task. However, these benchmarks do not focus on the more challenging aspect of cross-task zero-shot evaluation, where models must generalize to entirely new tasks with unseen combinations of object categories and motions. This gap in evaluation limits our understanding of the true generalization capabilities of VLA models. Motivated by this, this work proposes the first cross-task zero-shot manipulation benchmark, expanding the evaluation scope of the generalization capabilities of VLAs.
@@
In-context Learning with Large Language Models (LLMs). LLMs [65, 42, 66, 67, 68, 69, 70, 71] have demonstrated a remarkable capability known as in-context learning (ICL) [42, 43, 44], where they can learn to perform new tasks based on a few examples provided within the prompt, without requiring updates to parameters. The potential of ICL is being explored in robotics [72, 73, 74, 75, 76, 77]. Prior works like KAT [72] and RoboPrompt [76] have shown that off-the-shelf LLMs can predict robot actions directly using within-task in-context samples. By extending their paradigms to a cross-task setting, this work specifically focuses on the cross-task zero-shot generalization problem. Our X-ICM model uses demonstrations from seen tasks as in-context examples to prompt LLMs for action generation in completely unseen tasks. Under the cross-task zero-shot setting, the selection of in-context samples [78, 79, 80, 81] is crucial for a robust generalization. Our X-ICM method addresses this challenge by introducing a dynamics-guided sample selection strategy, ensuring that the selected seen demonstrations are relevant to the unseen tasks, thereby stimulating the cross-task generalization capabilities of LLMs.
3 AGNOSTOS: Cross-task Zero-shot Generalization Benchmark
@@
To systematically assess the cross-task zero-shot generalization ability of vision-language-action (VLA) models, we introduce AGNOSTOS, a reproducible benchmark built upon the RLBench simulation environment. AGNOSTOS features 18 seen tasks for training and 23 unseen tasks for rigorous generalization testing.
@@
Benchmark Simulator No. of train tasks No. of test tasks Cross-task Zero-shot Generalization Test
@@
Table 1 compares AGNOSTOS with existing robotic manipulation benchmarks in terms of their support for cross-task evaluation, i.e., considering test tasks that have different object categories or motions from training tasks. While many benchmarks focus on within-task visual generalization, few explicitly support cross-task zero-shot generalization test, and even fewer include tasks as challenging as our Level-2 testing scenarios. Even benchmarks that include some form of cross-task evaluation [^2] We carefully categorize their test tasks into our defined two-level difficulty task sets. often test on a narrow set of tasks, and typically evaluate only in-domain models, neglecting foundation and human-video pre-trained VLAs. AGNOSTOS fills this gap by supporting broad model coverage and introducing novel, difficult tasks in Level-2, providing a more comprehensive and diagnostic assessment of cross-task generalization.
@@ 4 X-ICM: Cross-task In-context Manipulation Method
To push the boundaries of cross-task zero-shot generalization in vision-language-action (VLA) models, we propose a method called Cross-task In-context Manipulation (X-ICM). Leveraging the cross-task generalization capabilities of LLMs, X-ICM utilizes demonstrations from seen tasks as in-context examples. The dynamic characteristics of these examples are used to prompt the LLM to predict action sequences for unseen tasks. In contrast to prior works [76, 72] that apply LLMs’ in-context learning to within-task generalization, our work is the first to extend this paradigm to a cross-task zero-shot setting. A central challenge in this setting is that the selection of in-context demonstrations significantly affects generalization performance. To address this, we design a dynamics-guided sample selection module that measures similarities between dynamic representations of seen and unseen tasks to guide the selection process, resulting in improved cross-task generalization.
@@ 4.1 Problem Definition and Method Overview
We tackle the problem of cross-task zero-shot robotic manipulation by exploiting the in-context learning ability of off-the-shelf LLMs. We assume access to a dataset of demonstrations from seen tasks, denoted as $\mathcal{D}^{s}=\{V_{i}^{s},A_{i}^{s},L_{i}^{s}\}_{i=1}^{N}$, where $N$ is the total number of seen demonstrations, $V^{s}_{i}=\{v^{s}_{i,1},\cdots,v^{s}_{i,T}\}$, $A^{s}_{i}=\{a^{s}_{i,1},\cdots,a^{s}_{i,T}\}$, and $L^{s}_{i}$ are the visual observation sequence, action sequence, and language description of the $i$-th seen demonstration ($T$ is the sequence length, which varies for each demonstration). For an unseen task with the given initial visual observation $v^{u}$, and language description $L^{u}$, our cross-task zero-shot manipulation method can be formulated as follows:
$A^{u}_{pred}=\{a^{u}_{1},\cdots,a_{t}^{u}\}=LLM(\mathcal{P}(\mathcal{F}^{sel}(\mathcal{D}^{s}),v^{u},L^{u})),$ (1)
@@ 4.2 Dynamics-guided Sample Selection module
$\min_{\mathcal{G}}\mathbb{E}_{i,z,\epsilon\sim\mathcal{N}(0,I)}\left[\left\|\epsilon-\epsilon_{\mathcal{G}}(v^{s}_{i,T,z},z,v^{s}_{i,1},L^{s}_{i})\right\|_{2}^{2}\right],$ (2)
@@ 4.3 Cross-task In-context Prediction module
For a tested unseen task, we use the obtained $K$ most-relevant seen demonstrations to construct the in-context prompt for cross-task action prediction with LLMs. Our Cross-task In-context Prediction module follows the practices in existing within-task in-context manipulation methods [76, 72], but we are the first to extend this paradigm to the cross-task zero-shot setting.
Textualizing key information. As shown at the bottom of Figure 2, for each selected seen demonstration, we extract the 3D positions of the centers of objects and the key-action sequence. For the 3D positions of objects’ centers, we use GroundingDINO [89] to detect the 2D positions of objects, and then use the depth, intrinsic, and extrinsic information of the camera to obtain the 3D coordinates of objects’ centers under the robot-base coordinate system. Additionally, we evenly divide the workspace into $100\times 100\times 100$ grids, and then normalize the 3D coordinates of the objects’ centers into the grid’s coordinate system. Thus, for the $m$-th object $O_{k}^{m}$ in the $k$-th seen demonstration, we textualize it as $O_{k}^{m}=$$objectname$: [$x$, $y$, $z$]” (e.g., “block: [52, 50, 19]”). For the key-action sequence, following [90], we select the actions when the gripper state changes or the joint velocities are near zero. In this way, the execution process of a demonstration can be determined by several key-actions. We use the 7-dimensional state of the end-effector to represent the robot action, i.e., the 3D positions, the roll-pitch-yaw angles, and the gripper’s open/close. The 3D positions of the end-effector are also normalized into the grid’s coordinate system of the workspace. And the angles of end-effector are discretized into 72 bins, with each occupying 5 degrees. Thus, for the $t$-th key-action $A_{k}^{t}$ in the $k$-th seen demonstration, we textualize it as $A_{k}^{t}=$ “[$x$, $y$, $z$, $roll$, $pitch$, $yaw$, $gripper$]” (e.g., “[53, 57, 17, 0, 36, 53, 0]”).
@@
We textualize all $K$ selected seen demonstrations in this manner, and concatenate them in descending order based on their dynamics similarity scores to the tested unseen task. For the tested unseen task, we textualize it using its language description and object context, i.e., “$[L^{unseen},\{O^{1},\cdots,O^{m},\cdots\}]\to$”. Finally, the cross-task in-context prompt is formed by concatenating the system prompt, the textualized all seen demonstrations, and the textualized tested unseen task (see the bottom of Figure 2 and an example in Figure A6 of the Appendix). This prompt will serve as the input of LLMs, which incentivizes the cross-task generalization capability of LLMs to predict the key-action sequence for the tested unseen task.
@@ 5.1 Benchmarking Vision-Language-Action Models
Table 2 presents the cross-task zero-shot performance of our X-ICM models on 23 unseen tasks from the benchmark, including 13 Level-1 and 10 Level-2 tasks. These tasks are designed to evaluate the generalization ability of VLA models under varying levels of difficulty. We compare our models against a diverse set of baselines, including models trained on in-domain RLBench data, human videos, and LLM/VLM-based foundations. Key observations include:
@@ 5.3 Real-world Experiments
To evaluate real-world applicability, we test X-ICM on five physical manipulation tasks using an xArm7 robot arm, DH-Robotics gripper, and a third-person Orbbec camera: put block into bin, push button, put bottle into box, stack blocks, and stack cups. We collect 20 demonstrations for each of the five tasks. During testing, we use the demonstrations from the other four tasks to construct the cross-task in-context prompt, enabling zero-shot generalization. The Qwen2.5-72B-Instruct LLM is used as the backbone and the number of seen demonstrations is set to $18$. Each task is executed 20 times, and we report average success rates. Results shown at the bottom of Figure 4 demonstrate the strong real-world cross-task zero-shot performance of our approach. We provide additional details in Section A3 of the Appendix, and showcase some testing videos that include both successful and failed cases in the Supplementary Materials.
Figure 4: Results of five real-world tasks. The tests are conducted in a cross-task zero-shot manner.

### v3

# Exploring the Limits of Vision-Language-Action Manipulation in Cross-task Generalization

@@ Abstract
Figure 1: The proposed AGNOSTOS benchmark evaluates zero-shot cross-task generalization through two difficulty levels. Level-1 testing involves 13 unseen tasks sharing partial similarity (objects or motions) with seen tasks. Level-2 testing has 10 unseen tasks from entirely novel scenarios, requiring stronger generalization capabilities. We systematically assess three broad categories of vision-language-action models, revealing critical limits in their ability to adapt to unseen tasks.
@@ 1 Introduction
Vision-Language-Action (VLA) models [3, 4, 5, 6, 2, 7, 8, 1, 9] have motivated a new era of robotic manipulation [10, 11, 12, 13, 14, 15, 16] by integrating visual perception, language understanding, and action generation. Through large-scale pre-training on diverse data, including human videos [17, 18, 19], real or simulated cross-embodiment robotic demonstrations [3, 20, 21], VLA models can effectively generalize across visual variations within known tasks (i.e., within-task generalization of seen tasks), such as handling objects in novel scenes or with altered properties. However, the true promise of VLA models lies in their capacity to generalize across tasks: to handle previously unseen combinations of objects, goals, and actions without prior exposure. This capability-zero-shot cross-task generalization-is essential for real-world deployment, where robots are expected to tackle novel tasks as they arise dynamically.
Despite the rapid progress in VLA research, most prior work [3, 4, 7, 8, 1] has focused on generalization testing in real-world environments. However, these evaluations are typically non-reproducible, and rarely target cross-task (i.e., unseen task) zero-shot generalization. Recently, many works [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] have been devoted to developing comprehensive simulated benchmarks, which could be utilized to evaluate the generalization of VLA models. While promising, these efforts mainly assess within-task generalization, leaving zero-shot cross-task generalization largely unexplored.
To address this critical gap, we present AGNOSTOS, a novel benchmark for evaluating zero-shot cross-task generalization in robotic manipulation. Built on RLBench [25], our benchmark comprises 23 unseen tasks that are carefully curated to differ from 18 commonly used seen training tasks [34, 35]. As shown in Figure 1, to probe different aspects of generalization, these unseen tasks are categorized into two difficulty levels:
@@
Motivated by the success of in-context learning [42, 43, 44] in large language models (LLMs), we propose a Cross-task In-context Manipulation (X-ICM) method to address the challenges of zero-shot cross-task generalization. X-ICM uses demonstrations from seen tasks as in-context examples to prompt LLMs to generate action plans for unseen tasks. A key challenge under this setup is selecting relevant demonstrations; irrelevant prompts fail to activate appropriate knowledge in the LLM, leading to poor cross-task predictions. To address this, we introduce a dynamics-guided sample selection strategy. This approach leverages learned representations of task dynamics—captured by predicting final observations from initial states and task descriptions—to identify relevant demonstrations across tasks. Guided by the learned dynamics, the dynamics-aware prompts can be formed that enable X-ICM to elicit high-quality action predictions from LLMs, even for unfamiliar, unseen tasks.
@@
  • We introduce AGNOSTOS, the first benchmark to systematically evaluate zero-shot cross-task generalization in robotic manipulation for VLA models, with 23 unseen tasks spanning two levels of generalization difficulty.
  • We propose X-ICM, a novel method that combines in-context prompting and dynamics-guided sample selection to enhance zero-shot cross-task transfer in VLA models.
@@ 2 Related Works
Benchmarks for Vison-Language-Action Models. Evaluating the generalization capabilities of VLA models requires robust and comprehensive benchmarks. While existing VLA models mainly focus on real-world testing, which suffers from reproducibility issues and rarely focuses systematically on zero-shot generalization to unseen tasks. Recently, many simulated benchmarks [22, 23, 24, 25, 26, 27, 28, 30, 31, 32, 33] have been proposed to offer a controlled environment for rigorous evaluation. Colosseum [31] enables evaluation of models across 14 axes of perturbations, including changes in color and texture, etc. GemBench [32] designs four levels of generalization, spanning novel placements, rigid and articulated objects, and complex long-horizon tasks. We find that these benchmarks have predominantly concentrated on evaluating visual variations within tasks, assessing how well models can generalize to new scenes or altered attributes of objects within the same task. However, these benchmarks do not focus on the more challenging aspect of zero-shot cross-task evaluation, where models must generalize to entirely new tasks with unseen combinations of object categories and motions. This gap in evaluation limits our understanding of the true generalization capabilities of VLA models. Motivated by this, this work proposes the first zero-shot cross-task manipulation benchmark, expanding the evaluation scope of the generalization capabilities of VLA models.
@@
In-context Learning with Large Language Models (LLMs). LLMs [65, 42, 66, 67, 68, 69, 70, 71] have demonstrated a remarkable capability known as in-context learning (ICL) [42, 43, 44], where they can learn to perform new tasks based on a few examples provided within the prompt, without requiring updates to parameters. The potential of ICL is being explored in robotics [72, 73, 74, 75, 76, 77]. Prior works like KAT [72] and RoboPrompt [76] have shown that off-the-shelf LLMs can predict robot actions directly using within-task in-context samples. By extending their paradigms to a cross-task setting, this work specifically focuses on the zero-shot cross-task generalization problem. Our X-ICM model uses demonstrations from seen tasks as in-context examples to prompt LLMs for action generation in completely unseen tasks. Under the zero-shot cross-task setting, the selection of in-context samples [78, 79, 80, 81] is crucial for robust generalization. Our X-ICM method addresses this challenge by introducing a dynamics-guided sample selection strategy, ensuring that the selected seen demonstrations are relevant to the unseen tasks, thereby stimulating the cross-task generalization capabilities of LLMs.
3 AGNOSTOS: Zero-shot Cross-task Generalization Benchmark
@@
To systematically assess the zero-shot cross-task generalization ability of vision-language-action (VLA) models, we introduce AGNOSTOS, a reproducible benchmark built upon the RLBench simulation environment. AGNOSTOS features 18 seen tasks for training and 23 unseen tasks for rigorous generalization testing.
@@
Benchmark Simulator
No. of
train
tasks
No. of
test
tasks
Cross-task Zero-shot Generalization Test
@@
Table 1 compares AGNOSTOS with existing robotic manipulation benchmarks in terms of their support for cross-task evaluation, i.e., considering test tasks that have different object categories or motions from training tasks. While many benchmarks focus on within-task visual generalization, few explicitly support zero-shot cross-task generalization test, and even fewer include tasks as challenging as our Level-2 testing scenarios. Even benchmarks that include some form of cross-task evaluation [^2] We carefully categorize their test tasks into our defined two-level difficulty task sets. often test on a narrow set of tasks, and typically evaluate only in-domain models, neglecting foundation and human-video pre-trained VLA models. AGNOSTOS fills this gap by supporting broad model coverage and introducing novel, difficult tasks in Level-2, providing a more comprehensive and diagnostic assessment of cross-task generalization.
@@ 4 X-ICM: Cross-task In-context Manipulation Method
To push the boundaries of zero-shot cross-task generalization in vision-language-action (VLA) models, we propose a method called Cross-task In-context Manipulation (X-ICM). Leveraging the cross-task generalization capabilities of LLMs, X-ICM utilizes demonstrations from seen tasks as in-context examples. The dynamic characteristics of these examples are used to prompt the LLM to predict action sequences for unseen tasks. In contrast to prior works [76, 72] that apply LLMs’ in-context learning to within-task generalization, our work is the first to extend this paradigm to a zero-shot cross-task setting. A central challenge in this setting is that the selection of in-context demonstrations significantly affects generalization performance. To address this, we design a dynamics-guided sample selection module that measures similarities between dynamic representations of seen and unseen tasks to guide the selection process, resulting in improved cross-task generalization.
@@ 4.1 Problem Definition and Method Overview
We tackle the problem of zero-shot cross-task robotic manipulation by exploiting the in-context learning ability of off-the-shelf LLMs. We assume access to a dataset of demonstrations from seen tasks, denoted as $\mathcal{D}^{s}=\{V_{i}^{s},A_{i}^{s},L_{i}^{s}\}_{i=1}^{N}$, where $N$ is the total number of seen demonstrations, $V^{s}_{i}=\{v^{s}_{i,1},\@cdots,v^{s}_{i,T}\}$, $A^{s}_{i}=\{a^{s}_{i,1},\@cdots,a^{s}_{i,T}\}$, and $L^{s}_{i}$ are the visual observation sequence, action sequence, and language description of the $i$-th seen demonstration ($T$ is the sequence length, which varies for each demonstration). For an unseen task with the given initial visual observation $v^{u}$, and language description $L^{u}$, our zero-shot cross-task manipulation method can be formulated as follows:
$A^{u}_{pred}=\{a^{u}_{1},\@cdots,a_{t}^{u}\}=LLM(\mathcal{P}(\mathcal{F}^{sel}(\mathcal{D}^{s}),v^{u},L^{u})),$ (1)
@@ 4.2 Dynamics-guided Sample Selection module
[…]
@@ 4.3 Cross-task In-context Prediction module
For a tested unseen task, we use the obtained $K$ most-relevant seen demonstrations to construct the in-context prompt for cross-task action prediction with LLMs. Our Cross-task In-context Prediction module follows the practices in existing within-task in-context manipulation methods [76, 72], but we are the first to extend this paradigm to the zero-shot cross-task setting.
Textualizing key information. As shown at the bottom of Figure 2, for each selected seen demonstration, we extract the 3D positions of the centers of objects and the key-action sequence. For the 3D positions of objects’ centers, we use GroundingDINO [89] to detect the 2D positions of objects, and then use the depth, intrinsic, and extrinsic information of the camera to obtain the 3D coordinates of objects’ centers under the robot-base coordinate system. Additionally, we evenly divide the workspace into $100\times 100\times 100$ grids, and then normalize the 3D coordinates of the objects’ centers into the grid’s coordinate system. Thus, for the $m$-th object $O_{k}^{m}$ in the $k$-th seen demonstration, we textualize it as $O_{k}^{m}=$objectnameobjectname: [$x$, $y$, zz]” (e.g., “block: [52, 50, 19]”). For the key-action sequence, following [90], we select the actions when the gripper state changes or the joint velocities are near zero. In this way, the execution process of a demonstration can be determined by several key-actions. We use the 7-dimensional state of the end-effector to represent the robot action, i.e., the 3D positions, the roll-pitch-yaw angles, and the gripper’s open/close. The 3D positions of the end-effector are also normalized into the grid’s coordinate system of the workspace. And the angles of end-effector are discretized into 72 bins, with each occupying 5 degrees. Thus, for the $t$-th key-action $A_{k}^{t}$ in the $k$-th seen demonstration, we textualize it as $A_{k}^{t}=$ “[$x$, $y$, $z$, rollroll, pitchpitch, yawyaw, grippergripper]” (e.g., “[53, 57, 17, 0, 36, 53, 0]”).
@@
We textualize all $K$ selected seen demonstrations in this manner, and concatenate them in descending order based on their dynamics similarity scores to the tested unseen task. For the tested unseen task, we textualize it using its language description and object context, i.e., “($L^{unseen},\{O^{1},\@cdots,O^{m},\@cdots\}$$\to$”. Finally, the cross-task in-context prompt is formed by concatenating the system prompt, the textualized all seen demonstrations, and the textualized tested unseen task (see the bottom of Figure 2 and an example in Figure A6 of the Appendix). This prompt will serve as the input of LLMs, which incentivizes the cross-task generalization capability of LLMs to predict the key-action sequence for the tested unseen task.
@@ 5.1 Benchmarking Vision-Language-Action Models
Table 2 presents the zero-shot cross-task performance of our X-ICM models on 23 unseen tasks from the benchmark, including 13 Level-1 and 10 Level-2 tasks. These tasks are designed to evaluate the generalization ability of VLA models under varying levels of difficulty. We compare our models against a diverse set of baselines, including models trained on in-domain RLBench data, human videos, and LLM/VLM-based foundations. Key observations include:
@@ 5.3 Real-world Experiments
To evaluate real-world applicability, we test X-ICM on five physical manipulation tasks using an xArm7 robot arm, DH-Robotics gripper, and a third-person Orbbec camera: put block into bin, push button, put bottle into box, stack blocks, and stack cups. We collect 20 demonstrations for each of the five tasks. During testing, we use the demonstrations from the other four tasks to construct the cross-task in-context prompt, enabling zero-shot generalization. The Qwen2.5-72B-Instruct LLM is used as the backbone and the number of seen demonstrations is set to $18$. Each task is executed 20 times, and we report average success rates. Results shown at the bottom of Figure 4 demonstrate the strong real-world zero-shot cross-task performance of our approach. We provide additional details in Section A3 of the Appendix, and showcase some testing videos that include both successful and failed cases in the Supplementary Materials.
Figure 4: Results of five real-world tasks. The tests are conducted in a zero-shot cross-task manner.
In addition, we also evaluate the proposed X-ICM model on long-horizon tasks under the zero-shot cross-task setting. We use tasks put block into bin, push button, put bottle into box, stack blocks, and pull out the middle block as seen tasks, where we have 10 demonstrations for each task. The long-horizon task we evaluated is clean the table, according the robot’s visual observation, we decompose the long-horizon task into three sequential sub-tasks, ie, stack cups, put the stacked cups into plate, place the mango on top of the stacked cups. We evaluated this task over 20 rollouts, with sub-task success contingent on the success of all prior sub-tasks. For our X-ICM method, the success rates on sub-tasks "stack cups", "put the stacked cups into plate", "place the mango on top of the stacked cups." are 40%, 25%, and 5%, respectively. Thus, the overall success rate for the long-horizon task is 5%. These results align with expectations, as the sequential nature of long-horizon tasks leads to cumulative failure. The zero-shot cross-task setting exacerbates these difficulties. These findings highlight the need for more advanced algorithms to handle long-horizon tasks in zero-shot scenarios.